OpenStudy (anonymous):

A self-employed contractor nearing retirement made two investments totaling 15,000. In one year, these investments yielded $1023 in simple interest. Part of the money was invested at 6% and the rest at 7.5%. How much was invested at each rate? show work

OpenStudy (gw2011):

Let x=investment at 6% Let y=investment at 7.5% x+y=15,000 0.06x+0.075y=1023 Multiply the first equation by 0.075 to eliminate y and find x 0.075x+0.075y=1125 0.06x+0.075y=1023 Subtract this equation from the first one to get: 0.015x=102 x=6800 Substitute this value into one of the equations--I used: x+y=15,000 y=15,000-6800=8200 Therefore, $6800 is invested at 6% and $8200 is invested at 7.5%